OpenStudy (anonymous):

In triangle ABC, a=6, b=9 and sin A=2/3. Find sin B

OpenStudy (anonymous):

wait so a and b are sides?

OpenStudy (anonymous):

In general if A is a corner, side a is the side opposite.

OpenStudy (anonymous):

but A, B, and C are probably the sides. a, b, and c are probably the angles

OpenStudy (anonymous):

If ABC is a triangle, they are the corners/ angles, ALWAYS.

OpenStudy (anonymous):

C /\ a / \ b /_____\ B c A

OpenStudy (anonymous):

C /\ 6 / \ 9 /_____\ B c sin(A) = 2/3 sin(B) = ?????

OpenStudy (anonymous):

sin A = a/c sin B = b/c sin A = 2/3 so a/c = 2/3 and a = 6 so 2/3 = 6/x x = 9 so side c = 9 sin B = b/c and b = 9 and c = 9 sin B = 1 I think.....

OpenStudy (anonymous):

thats the answer I got

OpenStudy (anonymous):

what about this problem: in triangle DEF angle D =40, DE=12 meters and DF=8 meters. Find the area of triangle DEF to the nearest tenth of a square meter